NEW CHIRPED-PULSE THZ FOURIER TRANSFORM TECHNIQIES FOR DETERMINATION OF LINE-SHAPE PARAMETERS FOR ATMOSPHERIC SPIECIES

EYAL GERECHT, \underline{KEVIN O. DOUGLASS}, DAVID F. PLUSQUELLIC, NATIONAL INSTITUTE OF STANDARDS AND TECHNOLOGY, OPTICAL TECHNOLOGY DIVISION, GAITHERS-BURG, MD 20899.

We will discuss the implementation of the recently developed Chirped-Pulse Fourier Transform THz spectrometer for the measurement of important atmospheric species. We will discuss how the method is used to obtain high-precision and high-sensitivity measurements of shapes, intensities, and broadening parameters directly from the rotational free induction decay signal. The current system measures a bandwidth of 10.6 GHz in a single measurement step with a resolution of 20 kHz and achieves high sensitivity in 30 seconds. Measurements of nitrous oxide, OCS, and other atmospheric species in the HITRAN database will be presented.